If solution A contains 30% salt and solution B contains 80% salt,a mixture needs 80 ounces that need to contain 65% salt, how much ounces would A and B both need each?
Answer by jim_thompson5910(35256)

Hint:

Translate the given statements/info into these equations

x+y = 80
0.3x+0.8y = 80*0.65

Solve the first equation for y to get y = 80 -x

Then plug this into the second equation to get

0.3x+0.8y = 80*0.65

0.3x+0.8(80 - x) = 80*0.65

Solve that for x, then use the solution for x to find the solution for y.

I'll let you do this. If you want to check the answers with me or if you still need help, see my email below. Thanks.


--------------------------------------------------------------------------------------------------------------

If you need more help, you can email me at jim_thompson5910@hotmail.com

To find other ways to contact me, you can also visit my website: http://www.freewebs.com/jimthompson5910/home.html

Thanks,

Jim

--------------------------------------------------------------------------------------------------------------